A sample of N2O3 contains 2.26 moles of nitrogen. Use the following atomic masses: N =14, O=16.

The number of moles of N2O3 present = mol N2O3
The mass of the sample = g N2O3
The number of moles of oxygen in the sample = mol O
The mass of oxygen in the sample = g O
Answer #1

(1)

Calculate the number of moles of oxygen in the sample:

\(\frac{2.26}{x}=\frac{2}{3} \Rightarrow x=\frac{2.26 \cdot 3}{2}=3.39\)

There are \(3.39\) moles of oxygen in the sample.

(2)

Calculate the mass of the oxgen in the sample:

\((16 \mathrm{~g} / \mathrm{mol})(3.39 \mathrm{~mol})=54.2 \mathrm{~g}\)

(3)

Calculate the mass of the nitrogen in the sample:

\((14 \mathrm{~g} / \mathrm{mol})(2.26 \mathrm{~mol})=31.6 \mathrm{~g}\)

(4)

Calculate the mass of the sample:

\(31.6 \mathrm{~g}+54.2 \mathrm{~g}=85.8 \mathrm{~g}\)

(5)

Calculate the molar mass of \(\mathrm{N}_{2} \mathrm{O}_{3}\) :

\((2 \mathrm{~mol})(14 \mathrm{~g} / \mathrm{mol})+(3 \mathrm{~mol})(16 \mathrm{~g} / \mathrm{mol})=76 \mathrm{~g} / \mathrm{mol}\)

(6)

Calulate the number of moles of \(\mathrm{N}_{2} \mathrm{O}_{3}\) present:

\(\frac{85.8 \mathrm{~g}}{76 \mathrm{~g} / \mathrm{mol}}=1.13 \mathrm{~mol}\)